### v1

@@ Abstract
Code cloning is frequently observed in software development, often leading to a variety of maintenance and security issues. While substantial research has been conducted on code cloning in traditional software, to the best of my knowledge, there is a lack of studies on cloning in VR software that consider its unique nature, particularly the presence of numerous serialized files in conjunction with the source code. In this paper, we conduct the first large-scale quantitative empirical analysis of software clones in 345 open-source VR projects, using the NiCad detector for source code clone detection and large language models (LLMs) for identifying serialized file clones. Our study leads to a number of insights into cloning phenomena in VR software, guided by seven carefully formulated research questions. These findings, along with their implications, are anticipated to provide useful guidance for both researchers and software developers within the VR field.
@@ 1 Introduction
In this study, we conduct the first quantitative empirical analysis of code cloning in 345 open-source VR projects, conducting clone detection in both source code and serialized files. For the source code portion of VR software, we use the NiCad tool [18, 15] for analysis, while for serialization files— which are difficult to detect using traditional methods— we utilize Large Language Models (LLMs) for analysis. Through seven carefully formulated research questions (RQs) and a set of proposed metrics, our study uncovers a number of insights into cloning phenomena in VR software and highlights the unique challenges of VR software development, offering practical guidance for researchers, developers, and maintainers within the VR field. Some of our insights are highlighted as follows: (i) Both source code cloning and serialized file cloning in VR software show a consistent pattern, with larger projects generally exhibiting more clones. (ii) The main difference between code clones in VR software and regular software lies in the presence of serialized files, which significantly reduce the level of source code cloning in a project. (iii) C# is the most commonly used programming language in open-source VR projects, whereas C is more likely to cause source code cloning issues. (iv) Third-party libraries introduced in VR applications often contribute to source code cloning issues, sometimes constituting the majority of the cloning results. (v) Inter-version code cloning is more common than intra-version code cloning, introducing significant security risks due to the potential propagation of vulnerabilities across multiple versions of the software. (vi) The cloning behavior of asset files varies with complexity. Complex files like scenes have lower clone levels, simpler ones like materials have higher, and prefab files fluctuate with development style.
@@
  • We formulate a suite of metrics to quantify code clones from various angles, enabling a comprehensive quantitative evaluation when used together.
@@
  • We derive several insights from the experimental results that deepen our understanding of VR software, providing practical guidance for researchers, developers, and maintainers in the VR domain.
@@ 2.2.1 Source Code Clone
2.2.2 Serialized Asset Files Clone
@@
Among these methods, content-based methods are simple and easy to implement, but they require significant computational resources and time, limiting their applicability to small-scale texts. Fingerprint-based methods, while computationally efficient, are sensitive to local changes and noise in the text. Deep learning-based methods, on the other hand, require substantial data and time for training, but with the advent of large models, this limitation has been significantly alleviated. In most scenarios, deep learning-based methods outperform the other approaches. The latest studies indicate that GPT-4, owing to its powerful contextual understanding and generative capabilities, often surpasses traditional embedding-based methods and other generative models in text similarity detection tasks, particularly those requiring reasoning and handling contextual variations [61, 62]. Given these facts, we propose using GPT-4o to identify clones in serialized asset files.
@@ 3 Methodology
The research methodology overview is shown in Figure 1. We begin by searching for VR projects in GitHub and existing datasets. After selecting 658 projects, we filter them by programming language, reducing the dataset to 416 projects. We then manually inspect and deduplicate these projects, resulting in a final dataset of 345 unique projects. Next, we formulate a set of metrics for measuring VR software cloning from various perspectives. We then propose detection methods for both source code cloning and serialized asset file cloning. Afterward, we conduct code cloning analysis from three angles to answer seven carefully formulated research questions.
@@ 3.2 Code Clone Detection
In this section, we establish a set of metrics to quantify the degree of cloning in both source code and serialized asset files, and provide detailed approaches for detecting these two cloning targets.
3.2.1 Metrics for Measuring Code Clone
For source code clone, we focus on code clone at the function granularity. Let $F$ denote the universe of functions in a VR software, and let ${\sf Sim}(func_{i},func_{j})(i\neq j)$ denote the similarity between function $func_{i}\in F$ and function $func_{j}\in F$. We define the concepts of clone function and clone class below.
$\exists func_{j}\in F(i\neq j),{\sf Sim}(func_{i},func_{j})>\tau$ (1)
Grounded in the above definitions, we put forward four metrics to evaluate the degree of source code cloning from both absolute and relative perspectives. The metrics applied to assess the absolute degree of code cloning are as follows:
  • Number of Clone Functions ($\sf NCF$): This metric measures the total number of clone functions in the source code of a VR software, namely
    ${\sf NCF}=|\{func_{i}\in F:\exists func_{j}\in F(i\neq j),{\sf Sim}(f_{i},f_{j})>\tau\}|$ (3)
  • Number of Clone Classes ($\sf NCC$): This metric measures the number of clone classes in the source code of a VR software, namely
    ${\sf NCC}=|\{C_{k}\subseteq F:\forall func_{i},func_{j}\in C_{k}(i\neq j),{\sf Sim}(func_{i},func_{j})>\tau\}|$ (4)
  • Ratio of Clone Functions ($\sf rcf$): This metric reflects the share of clone functions in a VR software, which is defined as the ratio of the number of clone functions ($\sf NCF$) to the total number of functions ($|F|$) in the software, namley
    ${\sf rcf}={\sf NCF}/|F|$ (5)
  • Ratio of Clone Classes ($\sf rcc$): This metric reflects the share of clone classes in a VR software, which is defined as the ratio of the number of clone classes ($\sf NCC$) to the total number of functions ($|F|$) in the software, namely
    ${\sf rcc}={\sf NCC}/|F|$ (6)
For serialized asset files clone, we focus on data clone at the file granularity. Let $A$ denote the universe of asset files in a VR software, and let ${\sf Sim}^{\prime}(file_{i},file_{j})(i\neq j)$ denote the similarity between asset file $file_{i}\in A$ and asset file $file_{j}\in A$. We define the concepts of clone file and clone group as follows.
Drawing from the definitions outlined above, we introduce a set of metrics to evaluate the extent of serialized asset cloning from both absolute and relative perspectives. The metrics applied to assess the absolute degree of asset cloning are as follows:
  • Number of Clone Assets ($\sf NCA$): This metric measures the total number of clone files in the serialized asset files of a VR software, namely
    ${\sf NCA}=|\{file_{i}\in A:\exists file_{j}\in A(i\neq j),{\sf Sim}^{\prime}(file_{i},file_{j})>\tau^{\prime}|$ (9)
  • Number of Clone Groups ($\sf NCG$): This metric measures the number of clone groups in the serialized asset files of a VR software, namely
    ${\sf NCG}=|\{G_{k}\subseteq A:\forall file_{i},file_{j}\in G_{k}(i\neq j),{\sf Sim}^{\prime}(file_{i},file_{j})>\tau^{\prime}\}|$ (10)
  • Ratio of Clone Assets ($\sf rca$): This metric reflects the share of clone files in a VR software, which is defined as the ratio of the number of clone files ($\sf NCA$) to the total number of asset files ($|A|$) in the software, namely
    ${\sf rca}={\sf NCA}/|A|$ (11)
  • Ratio of Clone Groups ($\sf rcg$): This metric reflects the share of clone groups in a VR software, which is defined as the ratio of the number of clone groups ($\sf NCG$) to the total number of asset files ($|A|$) in the software, namely
    ${\sf rcg}={\sf NCG}/|A|$ (12)
  • Clone Index ($\sf CI$): This metric measures the overall similarity of asset files in a VR software, computed by dividing the sum of pairwise similarities by the total number of asset files, namely
@@ 3.2.2 Source Code Clone Detection
Step 1 of the CoT framework requires the large language model to evaluate whether a single file pair qualifies as a clone pair. To facilitate this, four prompt types are developed: simple binary analysis, structure analysis, semantic analysis, and similarity analysis. Details of these prompts are provided in Table 2. Experimental evaluation of these prompts yields several key observations. Simple prompt leads to lower precision, as the model faces challenges in detecting inter-document clones effectively. Structure-based prompt results in high false positive rates, implying that the model is overly responsive to structural similarities within the files. Similarly, semantic dimension prompt also yields high false positive rates, likely due to the simple semantic structure of serialized asset files compared to natural or programming languages, leading to erroneous predictions. In contrast, prompt from the similarity analysis, which enables the model to generate a similarity score for file pairs, demonstrates the best performance and is subsequently chosen as the basis for future detection prompt.
@@ 3.3 Research Questions Definition
The objective of this study is to systematically investigate the phenomenon of code clones in VR software, aiming to clarify their characteristics, root causes, and implications, and to produce actionable insights for both researchers and practitioners alike. To achieve this goal, we design 7 research questions (RQs) from three perspectives: general cloning in software, cloning specific to source code, and cloning specific to serialized asset files.
@@
  • RQ1: Which VR projects are subjected to heavy cloning? We begin by examining the impact of different detection methods on clone detection results. Following this, we explore the degree of cloning in different VR software projects, considering both the source code and serialized file levels. Addressing this question will shed light on the projects with the most significant cloning, thereby assisting researchers and developers in prioritizing their efforts.
  • RQ2: What are the main differences in code cloning between VR and non-VR software? We first examine the structural differences between VR and non-VR software, and subsequently investigate how these structural variations contribute to the differences in code cloning between the two types of software. Addressing this question will clarify whether the observed findings are specific to VR software or reflect more universal trends within the software development domain.
  • RQ3: How to identify the most appropriate metrics for VR cloning analysis? We offer an in-depth analysis of the physical interpretations and applicability of the proposed metrics for measuring cloning. By addressing this question, we aim to guide users in selecting the most suitable metrics for various purposes and contexts.
@@
  • RQ4: Which programming languages dominate VR software development and are more susceptible to code cloning? We analyze the programming languages commonly used in VR software development, the code cloning issues they induce, and the underlying causes, aiming to identify which languages are more prone to cloning and why. Addressing this question will enable us to identify which projects are more prone to code cloning and allow for reflection on how language-level interventions might resolve these cloning challenges.
  • RQ5: How do third-party libraries impact code cloning in VR software? We explore the third-party libraries frequently utilized in VR projects and analyze whether their usage leads to code cloning and the specific patterns of cloning that arise. Addressing this question will help us identify the root causes of code cloning, whether it stems from the use of problematic third-party libraries or from developer practices, thereby allowing us to formulate targeted solutions.
  • RQ6: How do inter-version and intra-version code cloning evolve across different versions of a VR project? We investigate the presence of code cloning both within the same software version and between consecutive versions, exploring the degree of cloning and particularly focusing on the differences in cloning across different granularities of version changes. Addressing this question allows us to understand the progression of code cloning across software versions and uncover hidden security issues within these changes.
@@
  • RQ7: How do cloning practices vary across different types of asset files in VR software? We analyze the cloning discrepancies of various types of serialized asset files across multiple distinct projects. Addressing this question provides a finer understanding of asset file cloning, allowing developers to implement tailored maintenance practices for different types of assets.

### v2

@@ Abstract
Code cloning is frequently observed in software development, often leading to a variety of maintenance and security issues. While substantial research has been conducted on code cloning in traditional software, to the best of my knowledge, there is a lack of studies on cloning in virtual reality (VR) software that consider its unique nature, particularly the presence of numerous serialized files in conjunction with the source code. In this paper, we conduct the first large-scale quantitative empirical analysis of software clones in 345 open-source VR projects, using the NiCad detector for source code clone detection and large language models (LLMs) for identifying serialized file clones. Our study leads to a number of insights into cloning phenomena in VR software, guided by seven carefully formulated research questions. These findings, along with their implications, are anticipated to provide useful guidance for both researchers and software developers within the VR field.
@@ 1 Introduction
In this study, we substantially extend our previous work[17] through three key advancements: (i) we pioneer the exploration of clone-related issues in serialized files of VR software, a distinctive feature of VR systems compared to conventional software, whereas the prior study only examined source code clones. This is, to the best of our knowledge, the first work to analyze cloning in both source code and serialized files within VR software; (ii) we restructure the research framework by refining the original research questions (RQs) and evaluation metrics, and by formulating new VR-oriented questions, with the goal of shedding light on the mechanisms underlying code cloning in VR software and how these differ fundamentally from those observed in non-VR contexts; (iii) we considerably enlarge the dataset from 83 to 345 open-source VR projects, enhancing the robustness and generalizability of the findings.
Our study uncovers a number of insights into cloning phenomena in VR software and highlights the unique challenges of VR software development, offering practical guidance for researchers, developers, and industry practitioner within the VR field. Some of our insights are highlighted as follows: (i) Both source code cloning and serialized file cloning in VR software show a consistent pattern, with larger projects generally exhibiting more clones. (ii) Code and file cloning issues are most prevalent in gaming and education software, with gaming being the most significantly impacted, highlighting the need for targeted attention in this domain. (iii) Unlike traditional software, most VR projects exhibit significantly lower source code cloning due to their heavy reliance on serialized asset files. (iv) While C# is the most commonly used programming language in open-source VR projects, C-based projects exhibit a higher proportion of code cloning. (v) The introduction of third-party libraries in VR applications often contributes to source code cloning issues, sometimes constituting the majority of the cloning results. (vi) Inter-version code cloning is more common than intra-version code cloning, introducing significant security risks due to the potential propagation of vulnerabilities across multiple versions of the software. (vii) The cloning behavior of asset files varies with complexity. Complex files like scenes have lower clone levels, simpler ones like materials have higher, and prefab files fluctuate with development style.
@@
  • We derive several insights from the experimental results that deepen our understanding of VR software, providing practical guidance for researchers, developers, and industry practitioner in the VR domain.
@@ 2.2.1 Source Code Clone
2.2.2 Serialized File Clone
@@
Among these methods, content-based methods are simple and easy to implement, but they require significant computational resources and time, limiting their applicability to small-scale texts. Fingerprint-based methods, while computationally efficient, are sensitive to local changes and noise in the text. Deep learning-based methods, on the other hand, require substantial data and time for training, but with the advent of large models, this limitation has been significantly alleviated. In most scenarios, deep learning-based methods outperform the other approaches.
The nature of serialized asset files—often semi-structured, heterogeneous, and less amenable to syntax-based analysis—makes LLMs more suitable for capturing semantic similarities in this context. The latest studies indicate that GPT-4o, owing to its powerful contextual understanding and generative capabilities, often surpasses traditional embedding-based methods and other generative models in text similarity detection tasks, particularly those requiring reasoning and handling contextual variations [61, 62]. Given these facts, we propose using GPT-4o to identify clones in serialized asset files.
@@ 3 Methodology
The research methodology overview is shown in Figure 1. We begin by searching for VR projects in GitHub and existing datasets. After selecting 658 projects, we filter them by programming language, reducing the dataset to 416 projects. We then manually inspect and deduplicate these projects, resulting in a final dataset of 345 unique projects. We then propose detection methods for both source code cloning and serialized asset file cloning. Next, we formulate a set of metrics for measuring VR software cloning from various perspectives. Afterward, we conduct code cloning analysis from three angles to answer seven carefully formulated research questions.
@@ 3.2 Code Clone Detection
In this section, we provide detailed approaches for detecting cloning in both source code and serialized asset files, and define a set of metrics to quantify cloning from multi-faceted perspectives.
@@ 3.2.1 Source Code Clone Detection
Unlike traditional source code clones, asset clones often contain structured metadata, binary-encoded content, and heterogeneous formats. These characteristics demand semantic-level interpretation beyond surface-level syntax, thereby increasing the complexity of clone detection. While previous studies have demonstrated GPT-4o as the most effective LLM for text similarity detection tasks, it remains necessary to assess whether this conclusion holds in the context of asset file clone detection. Given the absence of a standardized dataset for this specific task, we design a three-step validation procedure: (i) independently evaluating 200 randomly selected file pairs using three different detection models; (ii) adopting the majority opinion (i.e., when two models agree) as a provisional benchmark in cases of conflicting results; and (iii) performing manual verification through a consensus-based review involving three researchers. As shown in Table 2, the final performance metrics confirm that GPT-4o remains the most effective model for asset file clone detection in our evaluation.
Table 2: Comparison of File Clone Detection Performance for Different LLMs.
Model Precision(%) Recall (%) F1-score (%) Accuracy (%)
GPT-4o 98.0 97.5 97.7 98.0
Claude 3.5 95.0 96.0 95.5 95.5
Gemini 1.5 pro 94.5 93.0 93.7 94.0
Note: (i) The consistency rate of the three models: 92.5% (185/200); (ii) GPT-4o got 11 edge cases correct.
@@
Step 1 of the CoT framework requires the LLM to evaluate whether a single file pair qualifies as a clone pair. To enable this process, we design four distinct prompt types: simple binary analysis, structural analysis, semantic analysis, and similarity analyses. Details of each prompt are provided in Table 3. We assess the effectiveness of these prompts by testing them on 200 randomly selected file pairs. The final performance is characterized using four standard metrics: precision, recall, F1-score, and accuracy. These metrics are derived through majority voting (using GPT-4o, Claude 3.5, Gemini 1.5 Pro), followed by manual inspection, in accordance with the three-step validation framework previously applied in LLM comparisons.
Table 4: Comparison of File Clone Detection Performance for Different Prompts.
Prompt Precision(%) Recall (%) F1-score (%) Accuracy (%)
Simple 52.5 40.0 45.0 45.0
Structure 60.0 80.0 68.6 55.0
Semantic 70.0 60.0 64.6 65.0
Similarity 98.0 97.5 97.7 98.0
Table 4 summarizes the experimental evaluation of the prompts. The simple prompt leads to the lowest performance, with precision at 52.5%, recall at 40.0%, F1-score at 45.0%, and accuracy at 45.0%, as the model struggles to effectively detect inter-document clones. The structure prompt achieves a higher recall of 80.0% but is hindered by a high false positive rate, with precision at 60.0%. This suggests that the model is overly sensitive to structural similarities within the files. The semantic prompt demonstrates moderate performance, with precision at 70.0%, recall at 60.0%, F1-score at 64.6%, and accuracy at 65.0%. This may be due to the simplified structure of serialized asset files compared to natural language. In contrast, the similarity analysis prompt excels with near-perfect precision at 98%, recall at 97.5%, F1-score at 97.7%, and accuracy at 98.0%, validating its ability in generating reliable similarity scores for file pairs.
@@
3.2.3 Metrics for Measuring Code Clone
For source code clone, we focus on code clone at the function granularity. Let $F$ denote the universe of functions in a VR software, and let ${\sf Sim}(func_{i},func_{j})(i\neq j)$ denote the similarity between function $func_{i}\in F$ and function $func_{j}\in F$. We define the concepts of clone function and clone class below. It is important to note that the similarity between functions is computed using the NiCad tool, as described in Section 3.2.1.
$\exists func_{j}\in F(i\neq j),{\sf Sim}(func_{i},func_{j})>\tau$ (1)
Grounded in the above definitions, we employ four straightforward metrics to evaluate the degree of source code cloning from both absolute and relative perspectives, following the existing studies [71, 5, 23, 17]. The metrics applied to assess the absolute degree of code cloning are as follows:
  • Number of Clone Functions ($\sf NCF$): This metric measures the total number of clone functions in the source code of a VR software.
  • Number of Clone Classes ($\sf NCC$): This metric measures the number of clone classes in the source code of a VR software.
  • Ratio of Clone Functions ($\sf rcf$): This metric reflects the share of clone functions in a VR software, which is defined as the ratio of the number of clone functions ($\sf NCF$) to the total number of functions ($|F|$) in the software.
  • Ratio of Clone Classes ($\sf rcc$): This metric reflects the share of clone classes in a VR software, which is defined as the ratio of the number of clone classes ($\sf NCC$) to the total number of functions ($|F|$) in the software.
For serialized asset files clone, we focus on data clone at the file granularity. Let $A$ denote the universe of asset files in a VR software, and let ${\sf Sim}^{\prime}(file_{i},file_{j})(i\neq j)$ denote the similarity between asset file $file_{i}\in A$ and asset file $file_{j}\in A$. We define the concepts of clone file and clone group as follows. It is important to note that the similarity between files is computed using the GPT-4o, as described in Section 3.2.2.
Analogous to source code clone metrics, we introduce a set of metrics to evaluate the extent of serialized asset cloning from both absolute and relative perspectives. The metrics applied to assess the absolute degree of asset cloning are as follows:
  • Number of Clone Assets ($\sf NCA$): This metric measures the total number of clone files in the serialized asset files of a VR software.
  • Number of Clone Groups ($\sf NCG$): This metric measures the number of clone groups in the serialized asset files of a VR software.
  • Ratio of Clone Assets ($\sf rca$): This metric reflects the share of clone files in a VR software, which is defined as the ratio of the number of clone files ($\sf NCA$) to the total number of asset files ($|A|$) in the software.
  • Ratio of Clone Groups ($\sf rcg$): This metric reflects the share of clone groups in a VR software, which is defined as the ratio of the number of clone groups ($\sf NCG$) to the total number of asset files ($|A|$) in the software.
To sum up, the proposed metrics ($\sf NCF$, $\sf rcf$) and ($\sf NCC$, $\sf rcc$) are based on clone functions and clone classes, as defined in Equations (1) and (2), respectively, in the context of code clones; whereas ($\sf NCA$, $\sf rca$) and ($\sf NCG$, $\sf rcg$) are based on clone files and clone groups, as defined in Equations (3) and (4), respectively, in the context of file clones.
To evaluate the overall similarity of asset files in a VR software, we introduce the Clone Index ($\sf CI$), a metric calculated by dividing the sum of pairwise similarities by the total number of asset files, namely
@@ 3.3 Research Questions Definition
The objective of this study is to systematically investigate the phenomenon of code clones in VR software, with the aim of clarifying their characteristics, root causes, and implications. By doing so, we aim to produce actionable insights for researcher, developers, and industry practitioners, ultimately improving the quality, maintainability, and performance of VR software. To achieve this goal, we propose seven research questions (RQs) from three analytical perspectives: (i) general cloning in software, (ii) cloning specific to source code, and (iii) cloning specific to serialized asset files.
@@
  • RQ1: Which VR projects are subjected to heavy cloning? The motivation behind RQ1 is to identify VR projects with significant levels of code and asset cloning, enabling developers to prioritize refactoring efforts and manage technical debt more effectively, providing researchers with insights into why certain projects accumulate more clones, and helping managers allocate resources toward maintaining high-risk systems. To answer this question, we first evaluate how various detection methods influence cloning results, and then assess the extent of cloning across VR projects at both the source code and asset file levels.
  • RQ2: What are the main differences in code cloning between VR and non-VR software? The motivation behind RQ2 is to understand how the unique structural and interactive features of VR software influence code cloning, since VR’s unique demands (e.g., real-time rendering, physics simulation, device I/O) may lead to distinct cloning patterns that differ from traditional software, thereby enabling developers to adopt best practices from non-VR domains where applicable, guiding researchers to design VR-specific clone detection tools if distinctive patterns are observed, and assisting software architects in deciding whether to invest in dedicated infrastructure for VR development. To answer this question, we first analyze the structural distinctions between VR and non-VR software, then assess how these differences contribute to variations in code cloning.
  • RQ3: How to measure and analyze code cloning in VR software development? The motivation behind RQ3 is to explore and evaluate the suitability of various cloning metrics in the context of VR software, as the unique characteristics of VR systems may require domain-specific measurement approaches, enabling developers to accurately assess code and asset maintainability, and helping researchers and tool builders establish standardized, context-aware evaluation criteria for clone analysis. To answer this question, we offer an in-depth analysis of the physical interpretations and applicability of the proposed metrics for measuring cloning.
@@
  • RQ4: Which programming languages dominate VR software development and are more susceptible to code cloning? The motivation behind RQ4 is to investigate how the choice of programming languages in VR development influences code cloning, since some languages may inherently promote more duplication than others, enabling developers to mitigate language-specific risks through targeted practices and guiding researchers in creating language-aware clone detection methods. To answer this question, we begin by identifying the dominant programming languages in VR development, then investigate how they contribute to code cloning and uncover the root causes behind these patterns.
  • RQ5: How do third-party libraries impact code cloning in VR software? The motivation behind RQ5 is to examine whether third-party libraries in VR development lead to specific cloning behaviors due to boilerplate or API constraints, helping developers identify dependency-related duplication risks, enabling researchers to distinguish between library-induced and developer-induced cloning patterns for more accurate detection, and informing library authors on improving API design to minimize clone proliferation. To answer this question, we explore the third-party libraries frequently utilized in VR projects and analyze whether their usage leads to code cloning and the specific patterns of cloning that arise.
  • RQ6: How do intra-version and inter-version code cloning evolve across different versions of a VR project? The motivation behind RQ6 is to examine the temporal dynamics of code cloning in VR projects, as understanding when and why duplication accumulates can help developers schedule refactoring milestones and assist researchers in identifying patterns that inform version-aware maintenance. To answer this question, we investigate the presence of code cloning both within the same software version and between consecutive versions, exploring the degree of cloning and particularly focusing on the differences in cloning across different granularities of version changes.
@@
  • RQ7: How do cloning practices vary across different types of asset files in VR software? The motivation behind RQ7 is to investigate cloning differences among various asset types in VR software, since VR-specific asset types (e.g., 3D models, scenes, shaders) may exhibit different cloning characteristics than traditional software artifacts, helping developers implement asset-specific maintenance strategies, aiding researchers in expanding non-code clone detection in content-rich environments, and assisting asset managers in refining asset control workflows. To answer this question, we analyze the cloning discrepancies of various types of serialized asset files across multiple distinct projects.